# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v1

@@ 2.1 Model and ex–ante objective in estimated diffusion markets
Fix $\theta\in\Theta$ and suppose, for the moment, that $\theta$ were observable to the controller. In Markovian settings with sufficient smoothness, the $\theta$-conditional value function $V^{\star,\theta}(t,x,y)$ satisfies an HJB equation whose control Hamiltonian (the part depending on $\pi$) can be written explicitly using (8):
$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,V_{x},V_{xx},V_{xy}):=x\,\pi^{\top}b(y,\theta)\,V_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,V_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,V_{xy},$ (15)
where $V_{x},V_{xx}$ are evaluated at $(t,x,y)$ and $V_{xy}(t,x,y)\in\mathbb{R}^{m}$. The last term in (15) is the return–factor hedging term induced by $d\langle W,W^{Y}\rangle\neq 0$.
The pointwise first-order condition for an interior optimizer is
$\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}=x\,V_{x}^{\star,\theta}\,b(y,\theta)+x^{2}\,V_{xx}^{\star,\theta}\,\Sigma(y,\theta)\,\pi+x\,\Sigma_{SY}(y,\theta)\,V_{xy}^{\star,\theta}\;=\;0.$ (16)
Assuming $\Sigma(y,\theta)$ is invertible and $V_{xx}^{\star,\theta}<0$, this yields the closed-form $\theta$-conditional full-information portfolio rule
@@
where $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ denotes the $\theta$-conditional Pontryagin sensitivity objects associated with the fixed policy $\pi$ in the frozen-$\theta$ market. Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
@@
The conditional first-variation identity (18) is standard for diffusion control under a fixed parameter $\theta$ (e.g. yong1999stochastic; fleming2006controlled; pham2009continuous). Taking the outer expectation over $\theta\sim q$ yields (19). Since $h$ is an arbitrary $\theta$-blind admissible perturbation, vanishing of the first variation at an interior optimum implies (20). For the quadratic portfolio Hamiltonian (15), substituting the explicit expression for $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ and introducing (21)–(23) yields the linear system (24) and the projected form (25) whenever $A_{t}$ is invertible. ∎
@@
If one allows history-dependent policies that explicitly infer $\theta$ from observed returns, a principled partial-information formulation introduces a time-varying posterior/belief state $q_{t}(\cdot)=\mathbb{P}(\theta\in\cdot\mid\mathcal{F}_{t}^{\mathrm{obs}})$. In such belief-state problems, the corresponding PMP/Hamiltonian criticality condition is expressed in terms of conditional expectations under $q_{t}$ (or, equivalently, conditional on $\mathcal{F}_{t}^{\mathrm{obs}}$); see, e.g., haussmann1987maximum; li1995general; baghery2007maximum. We do not pursue that learning/belief-state route here. Our algorithms and theory target the fixed-$q$, $q$-aggregated projection (25) under the $\theta$-blind Markov feedback restriction (52).
@@ 2.3 Gaussian references at a fixed decision time
This subsection collects Gaussian benchmarks that isolate decision-time statistical uncertainty and yield closed-form reference allocations. We fix a calendar decision time $t_{0}$ at which an external estimation pipeline outputs an uncertainty law $q_{t_{0}}(d\theta)$ for a risk-premium parameter, and we treat this law as an $\mathcal{F}_{t_{0}}$-measurable input for portfolio choice over the remaining horizon. This interface accommodates both Bayesian posterior/prior-like uncertainty descriptions (e.g., barberis2000investor; pastor2000portfolio) and frequentist sampling/resampling laws conditional on $\mathcal{F}_{t_{0}}$ (e.g., bootstrap or bagging) (e.g., efron1979bootstrap; efron1994bootstrap; breiman1996bagging). Throughout this subsection we work conditionally on $\mathcal{F}_{t_{0}}$, suppress conditioning by writing $q(d\theta)$, and shift the trading clock so that the decision time becomes 0 and the remaining horizon is $T$. These references are used as analytic targets and sanity checks for our numerical sections (Sections 4 and 5), rather than as characterizations of the unrestricted optimum of (14) over the full feedback class.
We present two decision-time references. Section 2.3.1 considers static drift uncertainty: a latent premium is drawn from $q$ once at time 0 and kept fixed on $[0,T]$, providing the controlled benchmark used in the high-dimensional scaling/geometry experiments of Section 4. Section 2.3.2 considers a mean-reverting (OU) premium with Gaussian initial uncertainty, which induces a horizon-dependent Gaussian law for the time-averaged premium and yields a tractable closed-form reference used in the hedging-demand recovery study of Section 5. For completeness, an online linear–Gaussian illustration that produces a time-varying uncertainty law $q_{t}$ via Kalman–Bucy filtering is deferred to Appendix A; it is included only to motivate a plug-in (receding-horizon) decision-time workflow in which $q_{t}$ is treated as an externally updated input at each decision time, rather than solving the fully optimal belief-state control problem (e.g., bensoussan1985optimal; pham2017dynamic).
@@ 2.3.1 Static Gaussian drift uncertainty
We start from a time-homogeneous Gaussian benchmark in which the (vector) risk premium is an unobserved static parameter drawn at the decision time. The agent commits to a single $\theta$-blind policy, and all ex–ante uncertainty is summarized by the decision-time law $q$.
@@
Let $d$ risky assets satisfy
$\frac{dS_{t}}{S_{t}}=r\,\mathbf{1}\,dt+\theta\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (26)
where $\Sigma\in\mathbb{R}^{d\times d}$ is symmetric positive definite and the latent excess-return vector is drawn at time 0 as
$\theta\sim q(d\theta).$ (27)
A $\theta$-blind portfolio fraction process $\pi_{t}\in\mathbb{R}^{d}$ generates wealth
$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\big(r+\pi_{t}^{\top}\theta\big)\,dt+\pi_{t}^{\top}\Sigma^{1/2}\,dW_{t},\hskip 18.49988ptX_{0}^{\pi}=x>0,$ (28)
and we evaluate the ex–ante objective
$J(\pi):=\mathbb{E}_{\theta\sim q}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta\big]\Big].$ (29)
For reference, under full information and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ ($\gamma>0,\gamma\neq 1$), the oracle Merton rule is $\pi^{\star}(\theta)=\frac{1}{\gamma}\Sigma^{-1}\theta$ (merton1969lifetime; merton1971optimum), which is infeasible here because $\theta$ is latent.
Analytic $q$-references via constant portfolios. To obtain a transparent closed-form benchmark that depends only on the decision-time law $q$, we temporarily restrict attention to constant portfolio fractions
$\pi_{t}\equiv\pi\in\mathbb{R}^{d}.$ (30)
This restriction is used solely to define an analytic $q$-reference; it is not imposed on the learning problem.
Log utility ($\gamma=1$).
Let $m_{\theta}:=\mathbb{E}_{\theta\sim q}[\theta]$. For constant $\pi$, the objective depends on $q$ only through $m_{\theta}$, and the optimal constant portfolio is
$\pi_{q,\log}^{\mathrm{const}}(T)=\Sigma^{-1}m_{\theta},$ (31)
which is independent of $T$ in this static benchmark. In the one-asset case ($d=1$, $\Sigma=\sigma^{2}$),
$\pi_{q,\log}^{\mathrm{const}}(T)=\frac{m_{\theta}}{\sigma^{2}}.$ (32)
CRRA ($\gamma\neq 1$): tilted optimality and Gaussian shrinkage.
For $\gamma\neq 1$ and constant $\pi$, conditional on $\theta$ the terminal wealth is lognormal, and
$J(\pi)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)rT-\tfrac{1}{2}\gamma(1-\gamma)T\,\pi^{\top}\Sigma\pi\Big\}\,M_{q}\big((1-\gamma)T\pi\big),$ (33)
where $M_{q}(u):=\mathbb{E}_{\theta\sim q}[\exp(u^{\top}\theta)]$ is the moment generating function of $q$. Any interior optimizer $\pi_{q,\gamma}^{\mathrm{const}}(T)$ satisfies the tilted first-order condition
$\gamma\,\Sigma\,\pi_{q,\gamma}^{\mathrm{const}}(T)=\nabla_{u}\log M_{q}(u)\Big|_{u=(1-\gamma)T\,\pi_{q,\gamma}^{\mathrm{const}}(T)}.$ (34)
If $q$ is Gaussian,
$\theta\sim\mathcal{N}(m_{\theta},P),\hskip 18.49988ptP\succeq 0,$ (35)
then $\nabla_{u}\log M_{q}(u)=m_{\theta}+Pu$ and the reference reduces to the linear system
$\big(\gamma\Sigma-(1-\gamma)T\,P\big)\,\pi_{q,\gamma}^{\mathrm{const}}(T)=m_{\theta},$ (36)
@@ A $\theta$-blind portfolio fraction process $\pi_{t}\in\mathbb{R}^{d}$ generates wealth
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\big(\gamma\Sigma-(1-\gamma)T\,P\big)^{-1}m_{\theta}.$ (37)
@@
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\big(\gamma\Sigma+(\gamma-1)T\,P\big)^{-1}m_{\theta},\hskip 18.49988pt(\gamma>1),$ (38)
and in one dimension ($d=1$, $\Sigma=\sigma^{2}$, $P=p$),
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\frac{m_{\theta}}{\gamma\sigma^{2}+(\gamma-1)T\,p},\hskip 18.49988pt(\gamma>1).$ (39)
@@ 2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
We next replace the static premium by a mean-reverting Gaussian premium process, a standard reduced-form device for return predictability and intertemporal hedging (campbell2002strategic; xia2001learning). Our goal here is not to introduce additional information structure, but to obtain a closed-form, decision-time Gaussian reference for the time-averaged premium over the remaining horizon. This induces a horizon-dependent effective premium law that can be used as a controlled analytic input in numerical experiments.
OU premium dynamics and decision-time uncertainty.
Let the uncertain initial state be $\vartheta$ and set $Y_{0}=\vartheta\sim\mathcal{N}(m_{0},P_{0})$, so the decision-time law is $q=\mathcal{N}(m_{0},P_{0})$. The premium factor follows
$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=\vartheta\sim\mathcal{N}(m_{0},P_{0}),$ (40)
and risky excess returns satisfy
$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},$ (41)
allowing instantaneous correlation
$d\langle W,W^{Y}\rangle_{t}=\rho\,dt,\hskip 18.49988pt\rho\in\mathbb{R}^{d\times m}.$ (42)
Integrated premium and induced Gaussian law.
Define the integrated premium
$I_{T}:=\int_{0}^{T}Y_{s}\,ds\in\mathbb{R}^{m}.$ (43)
Since (40) is linear-Gaussian with Gaussian initial condition, $I_{T}$ is Gaussian. Its mean and covariance are
$m_{I}(T)=\mathbb{E}[I_{T}]=T\bar{y}+K^{-1}\big(I-e^{-KT}\big)\,(m_{0}-\bar{y}),$ (44)
$C_{I}(T)=\mathrm{Cov}(I_{T})$ $=K^{-1}\big(I-e^{-KT}\big)\,P_{0}\,\big(I-e^{-KT}\big)^{\top}K^{-\top}$ (45)
$\qquad+\int_{0}^{T}K^{-1}\big(I-e^{-K(T-s)}\big)\,\Xi\Xi^{\top}\,\big(I-e^{-K(T-s)}\big)^{\top}K^{-\top}\,ds.$
Notably, (44)–(45) depend only on the OU dynamics and the decision-time uncertainty $(m_{0},P_{0})$; they do not depend on the return–factor correlation $\rho$.
Horizon-averaged premium and effective Gaussian law.
@@
Define the horizon-averaged effective premium
$\bar{\theta}_{T}:=\frac{1}{T}\,B\,I_{T}\in\mathbb{R}^{d}.$ (46)
Then
$\bar{\theta}_{T}\sim\mathcal{N}(m_{\bar{\theta}}(T),P_{\bar{\theta}}(T)),\hskip 18.49988ptm_{\bar{\theta}}(T)=\frac{1}{T}B\,m_{I}(T),\hskip 18.49988ptP_{\bar{\theta}}(T)=\frac{1}{T^{2}}B\,C_{I}(T)\,B^{\top}.$ (47)
When $\rho=0$, this induced law can be plugged directly into the static Gaussian reference of Section 2.3.1. When $\rho\neq 0$, the marginal law (47) remains valid, but constant-portfolio expected utility involves an additional cross-covariance term capturing the return–state shock linkage that generates hedging demand (campbell2002strategic; xia2001learning).
@@
Closed-form references under constant portfolios.
Restricting to constant fractions $\pi_{t}\equiv\pi$ turns the problem into a transparent decision-time benchmark: only the integrated premium $I_{T}=\int_{0}^{T}Y_{s}\,ds$ enters the drift of $\log X_{T}^{\pi}$, while the risk term remains time-homogeneous. This yields a closed-form target that depends on the decision-time law $q=\mathcal{N}(m_{0},P_{0})$ only through the induced mean $m_{I}(T)=\mathbb{E}[I_{T}]$ (and, for CRRA, through covariances as well).
Log utility ($\gamma=1$).
With $\pi_{t}\equiv\pi$, the log-utility criterion reduces to a strictly concave quadratic in $\pi$ whose drift term depends on the OU factor only through the mean integrated premium $m_{I}(T)=\mathbb{E}\!\left[\int_{0}^{T}Y_{s}\,ds\right]$. Hence the decision-time reference depends on $q=\mathcal{N}(m_{0},P_{0})$ only through $m_{\bar{\theta}}(T)=(1/T)B\,m_{I}(T)$ (and, in particular, does not involve return–factor correlation), giving
$\pi_{q,\log}^{\mathrm{const}}(T)=\Sigma^{-1}m_{\bar{\theta}}(T)=\frac{1}{T}\,\Sigma^{-1}B\,m_{I}(T).$ (48)
@@
Define
$C_{IW}(T):=\mathrm{Cov}(I_{T},W_{T})=\int_{0}^{T}K^{-1}\big(I-e^{-K(T-s)}\big)\,\Xi\,\rho^{\top}\,ds\;\in\;\mathbb{R}^{m\times d},$ (49)

### v2

@@ 2.1 Model and ex–ante objective in estimated diffusion markets
Fix $\theta\in\Theta$ and suppose, for the moment, that $\theta$ were observable to the controller. In Markovian settings with sufficient smoothness, the $\theta$-conditional value function $V^{\star,\theta}(t,x,y)$ satisfies an HJB equation whose control Hamiltonian (the part depending on $\pi$) can be written explicitly using (3). For clarity, we define the control Hamiltonian map as a function of generic “sensitivity” inputs $(\varphi_{x},\varphi_{xx},\varphi_{xy})$:
$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,\varphi_{x},\varphi_{xx},\varphi_{xy}):=x\,\pi^{\top}b(y,\theta)\,\varphi_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,\varphi_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,\varphi_{xy},$ (10)
where $\varphi_{x},\varphi_{xx}\in\mathbb{R}$ and $\varphi_{xy}\in\mathbb{R}^{m}$. In the full-information HJB interpretation, we evaluate (10) at $(\varphi_{x},\varphi_{xx},\varphi_{xy})=(V_{x},V_{xx},V_{xy})$. In the Pontryagin interpretation used below, the same affine map is evaluated at the corresponding $\theta$-conditional adjoint/sensitivity objects (denoted $(p^{\theta},p_{x}^{\theta},p_{y}^{\theta})$ along a trajectory); in smooth Markov regimes these objects coincide with the indicated value derivatives.
The pointwise first-order condition for an interior optimizer is therefore
$\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}=x\,V_{x}^{\star,\theta}\,b(y,\theta)+x^{2}\,V_{xx}^{\star,\theta}\,\Sigma(y,\theta)\,\pi+x\,\Sigma_{SY}(y,\theta)\,V_{xy}^{\star,\theta}\;=\;0,$ (11)
where $V_{x}^{\star,\theta},V_{xx}^{\star,\theta},V_{xy}^{\star,\theta}$ are evaluated at $(t,x,y)$. Assuming $\Sigma(y,\theta)$ is invertible and $V_{xx}^{\star,\theta}<0$, this yields the closed-form $\theta$-conditional full-information portfolio rule
@@
where $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ denotes the $\theta$-conditional Pontryagin sensitivity/adjoint objects associated with the fixed policy $\pi$ in the frozen-$\theta$ market. (Equivalently, (13) evaluates the same affine map $\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}$ defined by (10), with $(V_{x},V_{xx},V_{xy})$ replaced by the corresponding Pontryagin objects along the controlled trajectory.) Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
@@
The conditional first-variation identity (13) is standard for diffusion control under a fixed parameter $\theta$ (e.g. Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009). Taking the outer expectation over $\theta\sim q$ yields (14). Since $h$ is an arbitrary $\theta$-blind admissible perturbation, vanishing of the first variation at an interior optimum implies (15). For the quadratic portfolio Hamiltonian (10), substituting the explicit expression for $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ and introducing (16)–(18) yields the linear system (19) and the projected form (20) whenever $A_{t}$ is invertible. ∎
@@
If one allows history-dependent policies that explicitly infer $\theta$ from observed returns, a principled partial-information formulation introduces a time-varying posterior/belief state $q_{t}(\cdot)=\mathbb{P}(\theta\in\cdot\mid\mathcal{F}_{t}^{\mathrm{obs}})$. In such belief-state problems, the corresponding PMP/Hamiltonian criticality condition is expressed in terms of conditional expectations under $q_{t}$ (or, equivalently, conditional on $\mathcal{F}_{t}^{\mathrm{obs}}$); see, e.g., Haussmann (1987); Li and Tang (1995); Baghery and Øksendal (2007). We do not pursue that learning/belief-state route here. Our algorithms and theory target the fixed-$q$, $q$-aggregated projection (20) under the $\theta$-blind Markov feedback restriction.
@@ 2.3 Gaussian references at a fixed decision time
This subsection collects Gaussian benchmarks that yield closed-form decision-time reference allocations. Fix a decision time $t\in[0,T)$ and suppose an external procedure provides an $\mathcal{F}_{t}$-measurable input law $q_{t}$ for the uncertain premium parameter/state. We treat $q_{t}$ as a decision-time input (not a belief process updated within a run): it is held fixed over the remaining horizon $[t,T]$ and the resulting references are used only as validation targets and controlled diagnostics (not as characterizations of the unrestricted optimum). If the decision is revisited at a later time $t^{\prime}$, the reference can be recomputed by resolving the decision-time problem with the new input $q_{t^{\prime}}$ (plug-in replanning).
We present two decision-time references. Section 2.3.1 considers static drift uncertainty and provides the controlled benchmark used in the high-dimensional scaling/geometry experiments of Section 4. Section 2.3.2 considers a mean-reverting (OU) premium factor with Gaussian decision-time uncertainty for the current factor level, which induces a horizon-dependent effective Gaussian law for the time-averaged premium and yields a tractable reference used in the hedging-demand recovery study of Section 5. Appendix A gives a linear–Gaussian illustration of how an external filtering routine may produce a time-indexed input law $q_{t}$; it is included only as an input mechanism (plug-in replanning), not as a belief-state control formulation.
@@ 2.3.1 Static Gaussian drift uncertainty
We start from a time-homogeneous Gaussian benchmark in which the (vector) risk premium is an unobserved static parameter. This “unknown mean” specification is a standard device for isolating estimation risk/parameter uncertainty in portfolio choice (e.g., Kandel and Stambaugh, 1996; Barberis, 2000; Pástor, 2000). Fix a decision time $t\in[0,T)$ and remaining horizon $\tau:=T-t$. The agent chooses a $\theta$-blind policy on $[t,T]$, and all decision-time uncertainty about the (time-constant) premium is summarized by an $\mathcal{F}_{t}$-measurable law $q_{t}(d\theta)$, which we do not update over $[t,T]$.
@@
Let $d$ risky assets satisfy, for $s\in[t,T]$,
$\frac{dS_{s}}{S_{s}}=r\,\mathbf{1}\,ds+\theta\,ds+\Sigma^{1/2}dW_{s},\hskip 18.49988ptS_{t}\in(0,\infty)^{d},$ (21)
where $\Sigma\in\mathbb{R}^{d\times d}$ is symmetric positive definite and $\theta\in\mathbb{R}^{d}$ is a latent constant excess-return vector. At decision time $t$, uncertainty about $\theta$ is represented by the input law
$\theta\sim q_{t}(d\theta).$ (22)
A $\theta$-blind portfolio fraction process $\pi_{s}\in\mathbb{R}^{d}$ generates wealth on $[t,T]$:
$\frac{dX_{s}^{\pi}}{X_{s}^{\pi}}=\big(r+\pi_{s}^{\top}\theta\big)\,ds+\pi_{s}^{\top}\Sigma^{1/2}\,dW_{s},\hskip 18.49988ptX_{t}^{\pi}=x>0.$ (23)
We evaluate the (conditional) decision-time objective by mixing over $q_{t}$:
$J_{t}(\pi):=\mathbb{E}_{\theta\sim q_{t}}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta,\,\mathcal{F}_{t}\big]\Big].$ (24)
Under full information and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ ($\gamma>0,\gamma\neq 1$), the oracle Merton rule is $\pi^{\star}(\theta)=\frac{1}{\gamma}\Sigma^{-1}\theta$ (Merton, 1969, 1971), which is infeasible here because $\theta$ is latent.
Analytic $q_{t}$-references via portfolios constant over $[t,T]$. To obtain transparent closed-form targets, we temporarily restrict attention to portfolio fractions that are constant in trading time $s$ on $[t,T]$:
$\pi_{s}\equiv\pi\in\mathbb{R}^{d},\hskip 18.49988pts\in[t,T].$ (25)
The superscript $\mathrm{const}$” used below refers to this constancy in $s$. The resulting reference may still depend on the decision time $t$ through the input $q_{t}$ and on the remaining horizon $\tau=T-t$. (Decision-time constant-portfolio targets of this kind are closely connected to myopic/plug-in benchmarks used in the parameter-uncertainty literature; see, e.g., Barberis (2000); Pástor (2000).)
CRRA ($\gamma>1$): tilted optimality and Gaussian shrinkage.
For $\gamma>1$ and constant $\pi$, conditional on $\theta$ the terminal wealth is lognormal and
$J_{t}(\pi)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)r\tau-\tfrac{1}{2}\gamma(1-\gamma)\tau\,\pi^{\top}\Sigma\pi\Big\}\,M_{q_{t}}\big((1-\gamma)\tau\pi\big),$ (26)
where $M_{q_{t}}(u):=\mathbb{E}_{\theta\sim q_{t}}[\exp(u^{\top}\theta)]$ is the moment generating function of $q_{t}$. Any interior optimizer $\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)$ satisfies the tilted first-order condition
$\gamma\,\Sigma\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\nabla_{u}\log M_{q_{t}}(u)\Big|_{u=(1-\gamma)\tau\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)}.$ (27)
If the decision-time input is Gaussian,
$q_{t}=\mathcal{N}(m_{t},P_{t}),\hskip 18.49988ptP_{t}\succeq 0,$ (28)
then $\nabla_{u}\log M_{q_{t}}(u)=m_{t}+P_{t}u$ and the reference reduces to the linear system
$\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=m_{t},$ (29)
@@ A $\theta$-blind portfolio fraction process $\pi_{s}\in\mathbb{R}^{d}$ generates wealth on $[t,T]$:
$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)^{-1}m_{t}.$ (30)
@@
$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{t}\big)^{-1}m_{t},\hskip 18.49988pt(\gamma>1),$ (31)
which makes explicit how decision-time uncertainty $P_{t}$ dampens exposures as the remaining horizon grows (a qualitative effect widely emphasized in return-predictability/parameter-uncertainty studies; see, e.g., Barberis (2000); Pástor (2000); Kandel and Stambaugh (1996)).
@@ 2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
We next replace the static premium by a mean-reverting Gaussian premium factor, a standard reduced-form device for return predictability and intertemporal hedging (Campbell and Viceira, 2002; Xia, 2001). Fix a decision time $t\in[0,T)$ and remaining horizon $\tau:=T-t$. At time $t$ we take as input a Gaussian law for the current factor level $Y_{t}$ (rather than a point estimate), capturing decision-time statistical uncertainty (e.g., interval-type uncertainty quantification, resampling variability, or asymptotic approximations). Holding this input fixed over $[t,T]$, OU propagation and time-averaging induce a horizon-dependent effective premium law; when $\rho\neq 0$, return–factor shock correlation adds a further quadratic term relevant for hedging demand (Campbell and Viceira, 2002; Xia, 2001).
OU premium dynamics.
Let $Y_{s}\in\mathbb{R}^{m}$ be a premium factor following
$dY_{s}=K(\bar{y}-Y_{s})\,ds+\Xi\,dW_{s}^{Y},\hskip 18.49988ptK\in\mathbb{R}^{m\times m}\ \text{Hurwitz and invertible},$ (32)
where $\bar{y}\in\mathbb{R}^{m}$ and $\Xi\in\mathbb{R}^{m\times m}$. Risky excess returns satisfy
$dR_{s}:=\frac{dS_{s}}{S_{s}}-r\mathbf{1}\,ds=BY_{s}\,ds+\Sigma^{1/2}\,dW_{s},$ (33)
with instantaneous correlation
$d\langle W,W^{Y}\rangle_{s}=\rho\,ds,\hskip 18.49988pt\rho\in\mathbb{R}^{d\times m}.$ (34)
Decision-time input law for the factor.
At decision time $t$, we take as input an $\mathcal{F}_{t}$-measurable Gaussian law for the current factor level,
$Y_{t}\sim q_{t}(dy)=\mathcal{N}(m_{t},P_{t}),\hskip 18.49988ptP_{t}\succeq 0.$ (35)
A common special case is purely model-implied predictive propagation from an initial uncertainty $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$ without conditioning on observations:
$m_{t}=\bar{y}+e^{-Kt}(m_{0}-\bar{y}),$
$P_{t}=e^{-Kt}P_{0}e^{-K^{\top}t}+\int_{0}^{t}e^{-K(t-s)}\,\Xi\Xi^{\top}\,e^{-K^{\top}(t-s)}\,ds,$ (36)
which is an OU transition (prior-predictive) law and should not be interpreted as an information update. (References for return predictability and OU-style investment-opportunity dynamics include Campbell and Viceira (2002); Xia (2001).)
Induced Gaussian law for the time-averaged premium over $[t,T]$.
@@
Define the remaining-horizon integrated factor and the associated horizon-averaged effective premium:
$I_{t,T}:=\int_{t}^{T}Y_{s}\,ds\in\mathbb{R}^{m},\hskip 18.49988pt\bar{\theta}_{t,\tau}:=\frac{1}{\tau}\,B\,I_{t,T}\in\mathbb{R}^{d}.$ (37)
Let
$A(\tau):=K^{-1}\big(I-e^{-K\tau}\big)\in\mathbb{R}^{m\times m}.$
Under the linear–Gaussian OU model, $I_{t,T}$ is Gaussian under the input law (35):
$I_{t,T}\sim\mathcal{N}\big(m_{I}(t,\tau),C_{I}(t,\tau)\big),$
with
$m_{I}(t,\tau)=\mathbb{E}[I_{t,T}]=\tau\bar{y}+A(\tau)\,(m_{t}-\bar{y}),$ (38)
$C_{I}(t,\tau):=\mathrm{Cov}(I_{t,T})=A(\tau)\,P_{t}\,A(\tau)^{\top}+\int_{0}^{\tau}A(u)\,\Xi\Xi^{\top}\,A(u)^{\top}\,du.$ (39)
Consequently, the horizon-averaged effective premium is Gaussian,
$\bar{\theta}_{t,\tau}\sim\mathcal{N}\!\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big),\hskip 18.49988ptm_{\bar{\theta}}(t,\tau)=\frac{1}{\tau}B\,m_{I}(t,\tau),\hskip 18.49988ptP_{\bar{\theta}}(t,\tau)=\frac{1}{\tau^{2}}B\,C_{I}(t,\tau)\,B^{\top}.$ (40)
The pair $\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big)$ is the horizon-dependent effective Gaussian law that we use as a controlled analytic input.
@@
Closed-form decision-time references under portfolios constant over $[t,T]$.
As in Section 2.3.1, we define an analytic reference by restricting to portfolio fractions that are constant in trading time $s$ over $[t,T]$, i.e. $\pi_{s}\equiv\pi\in\mathbb{R}^{d}$. The superscript $\mathrm{const}$” below refers to this constancy in $s$; the resulting reference may still depend on the decision time $t$ through $(m_{t},P_{t})$ and on the remaining horizon $\tau$ through OU averaging.
@@
When $\rho\neq 0$, return shocks and factor shocks are contemporaneously correlated, producing an additional quadratic term in the constant-portfolio objective (hedging-demand channel; see, e.g., Campbell and Viceira (2002); Xia (2001)). Define the cross-covariance between $I_{t,T}$ and the return Brownian increment $W_{T}-W_{t}$:
$C_{IW}(\tau):=\mathrm{Cov}\!\big(I_{t,T},\,W_{T}-W_{t}\big)=\int_{0}^{\tau}A(u)\,\Xi\,\rho^{\top}\,du\;\in\;\mathbb{R}^{m\times d},$ (41)